Tom has an income of £1000 which he spends as follows:
80% goes on necessities and 20% on luxuries. If his income elasticity of demand for necessities is 0.5 and his income falls by 20 percent then:

A How much will Tom spend on necessities after the income fall? (4 marks)







B How much will Tom spend on luxuries after the income fall? (1 mark)




C Are necessities a normal or an inferior good? (1 mark)




D What is Tom’s income elasticity of demand for luxuries? (4 marks)
1
Expert's answer
2015-07-09T00:00:42-0400
A Tom will spend on necessities after the income fall 20*0.5 = 10 percent more, so he will spend 1000*0.8*(1 - 0.1) = £720

B Tom will spend on luxuries 1000 - 720 = £280 after the income fall.

C Necessities are normal good, as demand for them decreases with the decrease in income.

D Tom’s income elasticity of demand for luxuries is Eid = (280/200 - 1)/0.2 = 2.